Reports 1-1 of 1 Clear search Modify search
yohei.nishino - 17:03 Friday 13 May 2022 (20719) Print this report
Problems with OMC transmittance and Finesse

Aso, Michael, and Nishino,

After GW, we assembled DCPDs, QPDs, beam dumps, and cables. Then, We checked OMC transmissivity and finesse for the first time after GW and found two issues.

Low transmissivity and finesse

Transmissivity was ~ 30% and finesse was 462 (they are supposed to be 95% and 800) and ~50% of beam power was lost. Maybe during the assembly, the mirrors got contamination. We are cleaning them by first contact now.

Strange beam spot in transmittance light

We found a strange beam spot flickering outside the TEM00 beam when scanning the laser frequency. It doesn't seem like HOM, but actually coming from the OMC cavity (I was not able to take a picture of the flickering.) It is quite unreasonable. One possibility is dust on the mirror surface deflecting the laser path and making two resonant paths. Another possibility is beam clipping inside the OMC2 tombstone. To adjust the transmission path of the cavity, OMC2 is aligned slightly away from the designed position (about one beam aperture size away from the center). But we cannot see clipping clearly. 

Next Monday, we will remove first contact and check transmissivity and finesse again, then will align the mirrors again if necessary.


Images attached to this report
Search Help